Shift report 20000721 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Fri Jul 21 2000 - 08:14:24 EDT

  • Next message: http account: "Shift report 20000721 08:00-16:00"

    ------------------------------------------------------------
    Date: Fri Jul 21 08:14:24 2000, Shift: 00:00-08:00
    Supervisor: JHL, with: ZIS
    Log book#3 pages 37 - 38
    
    Shift Summary:
    ------------------------------------------------------------
    Machine: 
    It has been under the "restoring machine status".
    No beam except a few drops at the beginning of the shift.
    Since 6:00 a.m., "Quench link problems. Experts will
    be called." 
    
    Data: A large cosmic ray data set has been taken for the 
    Tile gain calibrations.  
    
    Detectors: Some problematic(?) channels:
    H1: (Top 17, 18) Bottom 33
    H2: Top 1,5,7,13 (15,25,26)
    TOFW: Top 15, 34-40 Bottom: 21, 33-41, 57, (68, 69) 80
    BB: Right 9 
    
    A question for the Si people: There are two sets of
    temperature distributions ("SiTemp"). Some read around 10-20, 
    and the other ones are around 200-300.  Are they normal?
    And it's almost impossible to read the numbers on "SiTemp".
    Something has to be done for a better reading.
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Fri Jul 21 2000 - 08:17:56 EDT